# Toricity in families of Fano varieties

### v1

@@ 3. Toric log Calabi–Yau pairs in families
    the restriction of $B$ to $B|_{X_{t_{0}}}$ induces a bijection between prime components,
@@
Let $T^{\prime}\to T$ be a finite Galois cover such that for the pullback $(X_{T^{\prime}},B_{T^{\prime}})$, the prime components of $B_{T^{\prime}}$ restrict to prime divisors on general fibers of $X_{T^{\prime}}\to T^{\prime}$. Let $(X^{\prime\prime},B^{\prime\prime})\to(X_{T^{\prime}},B_{T^{\prime}})$ be a $\mathbb{Q}$-factorial dlt modification, which exists by [32, Theorem 3.1]. We get a commutative diagram
@@ 4. Cluster type log Calabi–Yau pairs in families
In order to spread out the birational contraction $\phi_{t_{0}}$, we will need to extend certain divisors over a family. An important ingredient for this is the following lemma on invariance of sections for suitable dlt modifications of Fano type morphisms. This lemma extends [26, Lemma 2.7].
@@
${(Y,B_{Y})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi}$${(X,B)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${X^{\prime}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\psi}$$\scriptstyle{\phi}$${T.}$
@@
By Lemma 2.23, every exceptional non-terminal place of $(X,\Gamma_{\epsilon})$ is a log canonical place of $(X,B)$. Therefore, every non-terminal center of $(X,\Gamma_{\epsilon})$ is a stratum of $(X,\lceil D^{\prime}_{\epsilon}\rceil$). By the assumption on the strata of $(X,B)$, every proper stratum of $(X^{\prime},\lceil D^{\prime}_{\epsilon}\rceil)$ is horizontal over $T$ and restricts to a proper stratum on each closed fiber. Therefore, by successively performing blow-ups of the strata of $(X^{\prime},\lceil D^{\prime}_{\epsilon}\rceil)$, we may assume that $(X^{\prime},D^{\prime}_{\epsilon})$ is a terminal pair (see, e.g., [34, Proposition 2.36]).
Let $\delta>0$ be a rational number small enough so that $A_{\epsilon,\delta}\coloneqq\epsilon A+\delta D_{X}$ is ample and $(X^{\prime},D^{\prime}_{\epsilon,\delta}\coloneqq D^{\prime}_{\epsilon}+\delta\phi^{*}E_{Y})$ is terminal. By [25, Theorem 1.8], for every sufficiently divisible $m$ and for every closed point $t\in U$, the homomorphism
@@
[…]
@@
We first construct $U\rightarrow T$ and $(X_{U},B_{U})$ as in (1). Since each fiber of $X\rightarrow T$ is ${\mathbb{Q}}$-factorial and terminal, it has rational singularities [14] and is smooth in codimension $2$. Moreover, by Kawamata–Viehweg vanishing, every fiber has acyclic structure sheaf, so we may apply [17, Theorem 4.1] to find a dominant finite étale morphism $T^{\prime}\rightarrow T$ from a smooth variety such that the base change $f^{\prime}\colon X^{\prime}\coloneqq X_{T^{\prime}}\rightarrow T^{\prime}$ has the property that ${\rm Cl}(X^{\prime}/T^{\prime})$ maps surjectively to the class group of each closed fiber. Let $(X^{\prime},B^{\prime})$ (resp. $(X^{\prime},\Gamma^{\prime})$) be the log pull-back of $(X,B)$ (resp. $(X,\Gamma)$) to $X^{\prime}$. Note that $f^{\prime}$ is still a Fano type fibration, $(X^{\prime},B^{\prime})$ is log Calabi–Yau of index one over $T^{\prime}$, and $(X^{\prime},\Gamma^{\prime})$ is terminal and log Calabi–Yau over $T^{\prime}$. By [35, Lemma 30] applied to any dlt modification of $(X^{\prime},B^{\prime})$, we can find a dominant finite étale morphism $U\rightarrow T^{\prime}$ from a smooth variety such that if $f_{U}\colon(X_{U},B_{U})\rightarrow T^{\prime\prime\prime}$ is the base change, then the dual complex of $(X_{U},B_{U})$ restricts to the dual complex of the log fibers. In particular, each prime component of $B_{U}$ remains prime after restriction to any closed fiber. We let $(X_{U},\Gamma_{U})$ be the log pullback of $(X^{\prime},\Gamma^{\prime})$ to $X_{U}$. The morphism $f_{U}$ is Fano type, $(X_{U},B_{U})$ is log CY of index one over $U$, and $(X_{U},\Gamma_{U})$ is terminal and log CY over $U$. Further, $\operatorname{Cl}(X_{U}/U)$ maps surjectively to the class group of each closed fiber of $f_{U}$.
@@
Consider the commutative diagram provided by Proposition 4.4:
[…]
@@
Next, since $Y_{U_{0}}\dashrightarrow W$ is a birational contraction and contracts precisely $E_{Y,U_{0}}$, the negativity lemma implies that $(W,B_{W})$ is a log CY pair and $(Y_{U_{0}},B_{Y,U_{0}})\dashrightarrow(W,B_{W})$ is crepant birational. Furthermore, over each closed point $s\in U_{0}$, the same argument and conclusions hold for the log fibers. Over the point $s_{0}$, since $E_{s_{0}}$ is $\pi_{s_{0}}$-exceptional, the composition $(W_{s_{0}},B_{W,s_{0}})\dashrightarrow(Y_{s_{0}},B_{Y,s_{0}})\dashrightarrow({\mathbb{P}}^{d},\Sigma^{d})$ is a crepant birational contraction that only contracts log canonical places of $({\mathbb{P}}^{d},\Sigma^{d})$. So $(W_{s_{0}},B_{W,s_{0}})$ is a toric log CY pair by Lemma 2.15. Thus, we have a $\mathbb{Q}$-factorial pair $(W,B_{W})$ that is log CY of index one and satisfies the following conditions:
@@ 5. Proofs of main theorems
By Noetherian induction, it suffices to show that $T_{\rm toric}\cap U$ is constructible for some open dense affine subset $U\subseteq T$. By [26, Lemma 2.7], there is a dense open subset $U\subset T$ such that for $m$ divisible enough, the map $H^{0}(\mathcal{X},-mK_{\mathcal{X}})\rightarrow H^{0}(\mathcal{X}_{t},-mK_{\mathcal{X}_{t}})$ is surjective (and nonzero) for every closed point $t\in U$. Fix such an $m$, let
@@
The $n$-dimensional smooth Fano varieties form a bounded family by [33, Theorem 0.2]; that is, there exist finitely many surjective projective morphisms $g_{i}\colon\mathcal{Z}_{i}\to S_{i}$ of quasi-projective varieties such that any $n$-dimensional smooth Fano variety is isomorphic to a closed fiber of $g_{i}$ for some $i$ (see, e.g., [3, Section 2.8]). Furthermore, by standard arguments, one may replace each $S_{i}$ by a constructible subset to assume every closed fiber is an $n$-dimensional smooth Fano variety, i.e., that the $g_{i}$ are parametrizing families for $n$-dimensional smooth Fano varieties. Then the corollary is immediate from Theorem 1.2. ∎

### v2

@@ 3. Toric log Calabi–Yau pairs in families
    the restriction of $B$ to $B|_{X_{t_{0}}}$ induces a bijection between prime components of $B$ and $B|_{X_{t_{0}}}$,
@@
Let $T^{\prime}\to T$ be a finite Galois cover such that for the pullback $(X_{T^{\prime}},B_{T^{\prime}})$, the prime components of $B_{T^{\prime}}$ restrict to prime divisors on general fibers of $X_{T^{\prime}}\to T^{\prime}$. Let $(X^{\prime\prime},B^{\prime\prime})\to(X_{T^{\prime}},B_{T^{\prime}})$ be a $\mathbb{Q}$-factorial dlt modification, which exists by [34, Theorem 3.1]. We get a commutative diagram
@@
We finish this section by proving a more general version of Theorem 1.1. Even though Theorem 3.1 requires a Fano type assumption, we retain the theorem statement and the above proof because it will be useful in proving Theorem 1.2. The proof of the following theorem was kindly suggested to us by János Kollár and Burt Totaro.
Theorem 3.4.
Let $\mathcal{X}\rightarrow T$ be a family of varieties (over $\mathbb{C}$). Then, the set
is a constructible subset of $T$.
Proof.
First, we assume that the closed fibers of $\mathcal{X}\rightarrow T$ are smooth varieties. Then, there is a dense open subset $U\subset T$ such that $U$ is smooth and the morphism $\mathcal{X}_{U}\to U$ is smooth, so Ehresmann’s theorem implies the closed fibers $\mathcal{X}_{t}$ have constant Betti numbers over $U$. We may assume that the first Betti number of each fiber over $U$ is 0, otherwise none of the fibers are toric (see, e.g., [9, Theorem 12.1.10]). By upper semicontinuity of $h^{0}(\mathcal{X},\mathcal{T}_{\mathcal{X}_{t}})$, we may restrict to a dense open subset $V\subset U$ over which $h^{0}(\mathcal{X},\mathcal{T}_{\mathcal{X}_{t}})$ is constant. Then $\operatorname{Aut}^{0}(\mathcal{X}_{V}/V)\to V$ is a smooth family of algebraic groups. Using the assumption that $b_{1}(\mathcal{X}_{t})=0$, we may apply [23, Théorème 1.7(a)] to conclude that the rank of $\operatorname{Aut}^{0}(\mathcal{X}_{t})$ is lower semicontinuous over $V$. Thus, we may further restrict to a dense open $W\subset V$ over which $\operatorname{Aut}^{0}(\mathcal{X}_{t})$ has constant rank. Over $W$, either all or none of the fibers are toric, since $\mathcal{X}_{t}$ is toric if and only if the rank of $\operatorname{Aut}^{0}(\mathcal{X}_{t})$ is equal to $\dim\mathcal{X}_{t}$. Then, Noetherian induction completes the proof in the case that the closed fibers of $\mathcal{X}\rightarrow T$ are smooth varieties.
Now, we let $\mathcal{X}\rightarrow T$ be any family of varieties. Let $\mathcal{Y}\rightarrow\mathcal{X}$ be a functorial resolution of singularities (see, e.g., [33, Theorem 3.36 and Theorem 3.35(5)]). Then, for $t\in T$ general, we have an induced functorial resolution $\mathcal{Y}_{t}\rightarrow\mathcal{X}_{t}$ (see, e.g., [33, 3.34.4]). So there is a dense open subset $U$ such that $\mathcal{Y}_{t}\rightarrow\mathcal{X}_{t}$ is a functorial resolution of singularities for each closed point $t\in U$. Therefore, the closed fibers of the morphism $\mathcal{Y}|_{U}\rightarrow U$ are smooth varieties. By the first paragraph, we conclude that
$U_{\rm toric}(\mathcal{Y})\coloneqq\{\text{closed $t\in U$}\mid\mathcal{Y}_{t}\text{ is a toric variety}\}$
is a constructible subset of $U$. Functorial resolutions are equivariant by [33, Proposition 3.9.1]; therefore, for $t\in U$, $\mathcal{Y}_{t}$ is toric whenever $\mathcal{X}_{t}$ is toric. On the other hand, image of toric varieties via projective birational morphism are also toric by Lemma 2.15. Therefore, for $t\in U$, $\mathcal{X}_{t}$ is toric whenever $\mathcal{Y}_{t}$ is toric. We conclude that $U_{\rm toric}(\mathcal{Y})=T_{\rm toric}\cap U$. Then, Noetherian induction completes the proof. ∎
@@ 4. Cluster type log Calabi–Yau pairs in families
In order to spread out the birational contraction $\phi_{t_{0}}$, we will need to extend certain divisors over a family. An important ingredient for this is the following lemma on invariance of sections for suitable dlt modifications of Fano type morphisms. This lemma extends [27, Lemma 2.7].
@@
[…]
@@
By Lemma 2.24, every exceptional non-terminal place of $(X,\Gamma_{\epsilon})$ is a log canonical place of $(X,B)$. Therefore, every non-terminal center of $(X,\Gamma_{\epsilon})$ is a stratum of $(X,\lceil D^{\prime}_{\epsilon}\rceil$). By the assumption on the strata of $(X,B)$, every proper stratum of $(X^{\prime},\lceil D^{\prime}_{\epsilon}\rceil)$ is horizontal over $T$ and restricts to a proper stratum on each closed fiber. Therefore, by successively performing blow-ups of the strata of $(X^{\prime},\lceil D^{\prime}_{\epsilon}\rceil)$, we may assume that $(X^{\prime},D^{\prime}_{\epsilon})$ is a terminal pair (see, e.g., [36, Proposition 2.36]).
Let $\delta>0$ be a rational number small enough so that $A_{\epsilon,\delta}\coloneqq\epsilon A+\delta D_{X}$ is ample and $(X^{\prime},D^{\prime}_{\epsilon,\delta}\coloneqq D^{\prime}_{\epsilon}+\delta\phi^{*}E_{Y})$ is terminal. By [26, Theorem 1.8], for every sufficiently divisible $m$ and for every closed point $t\in U$, the homomorphism
@@
${(X,B)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f}$${(X_{U},B_{U})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f_{U}}$$\scriptstyle{\phi}$${(Z,B_{Z})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f_{Z}}$${T}$${U\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
@@
We first construct $U\rightarrow T$ and $(X_{U},B_{U})$ as in (1). Since each fiber of $X\rightarrow T$ is $\mathbb{Q}$}-factorial and terminal, it has rational singularities [14] and is smooth in codimension two. Moreover, by Kawamata–Viehweg vanishing, every fiber has acyclic structure sheaf, so we may apply [17, Theorem 4.1] to find a dominant finite étale morphism $T^{\prime}\rightarrow T$ from a smooth variety such that the base change $f^{\prime}\colon X^{\prime}\coloneqq X_{T^{\prime}}\rightarrow T^{\prime}$ has the property that ${\rm Cl}(X^{\prime}/T^{\prime})$ maps surjectively to the class group of each closed fiber. Let $(X^{\prime},B^{\prime})$ (resp. $(X^{\prime},\Gamma^{\prime})$) be the log pull-back of $(X,B)$ (resp. $(X,\Gamma)$) to $X^{\prime}$. Note that $f^{\prime}$ is still a Fano type fibration, $(X^{\prime},B^{\prime})$ is log Calabi–Yau of index one over $T^{\prime}$, and $(X^{\prime},\Gamma^{\prime})$ is terminal and log Calabi–Yau over $T^{\prime}$. By [37, Lemma 30] applied to any dlt modification of $(X^{\prime},B^{\prime})$, we can find a dominant finite étale morphism $U\rightarrow T^{\prime}$ from a smooth variety such that if $f_{U}\colon(X_{U},B_{U})\rightarrow T^{\prime\prime\prime}$ is the base change, then the dual complex of $(X_{U},B_{U})$ restricts to the dual complex of the log fibers. In particular, each prime component of $B_{U}$ remains prime after restriction to any closed fiber. We let $(X_{U},\Gamma_{U})$ be the log pullback of $(X^{\prime},\Gamma^{\prime})$ to $X_{U}$. The morphism $f_{U}$ is Fano type, $(X_{U},B_{U})$ is log CY of index one over $U$, and $(X_{U},\Gamma_{U})$ is terminal and log CY over $U$. Further, $\operatorname{Cl}(X_{U}/U)$ maps surjectively to the class group of each closed fiber of $f_{U}$.
@@
Consider the following commutative diagram provided by Proposition 4.4.
${(X,B)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f}$${(X_{U},B_{U})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f_{U}}$$\scriptstyle{\phi}$${(Z,B_{Z})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f_{Z}}$${T}$${U\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
@@
Next, since $Y_{U_{0}}\dashrightarrow W$ is a birational contraction and contracts precisely $E_{Y,U_{0}}$, the negativity lemma implies that $(W,B_{W})$ is a relative log CY pair over $U$ and $(Y_{U_{0}},B_{Y,U_{0}})\dashrightarrow(W,B_{W})$ is relative crepant birational over $U$. Furthermore, over each closed point $s\in U_{0}$, the same argument and conclusions hold for the log fibers. Over the point $s_{0}$, since $E_{s_{0}}$ is $\pi_{s_{0}}$-exceptional, the composition $(W_{s_{0}},B_{W,s_{0}})\dashrightarrow(Y_{s_{0}},B_{Y,s_{0}})\dashrightarrow({\mathbb{P}}^{d},\Sigma^{d})$ is a crepant birational contraction that only contracts log canonical places of $({\mathbb{P}}^{d},\Sigma^{d})$. So $(W_{s_{0}},B_{W,s_{0}})$ is a toric log CY pair by Lemma 2.16. Thus, we have a $\mathbb{Q}$-factorial pair $(W,B_{W})$ that is log CY of index one and satisfies the following conditions:
@@ 5. Proofs of main theorems
By Noetherian induction, it suffices to show that $T_{\rm toric}\cap U$ is constructible for some open dense affine subset $U\subseteq T$. By [27, Lemma 2.7], there is a dense open subset $U\subset T$ such that for $m$ divisible enough, the map $H^{0}(\mathcal{X},-mK_{\mathcal{X}})\rightarrow H^{0}(\mathcal{X}_{t},-mK_{\mathcal{X}_{t}})$ is surjective (and nonzero) for every closed point $t\in U$. Fix such an $m$, let
@@
The $n$-dimensional smooth Fano varieties form a bounded family by [35, Theorem 0.2]; that is, there exist finitely many surjective projective morphisms $g_{i}\colon\mathcal{Z}_{i}\to S_{i}$ of quasi-projective varieties such that any $n$-dimensional smooth Fano variety is isomorphic to a closed fiber of $g_{i}$ for some $i$ (see, e.g., [3, Section 2.8]). Furthermore, by standard arguments, one may replace each $S_{i}$ by a constructible subset to assume every closed fiber is an $n$-dimensional smooth Fano variety, i.e., that the $g_{i}$ are parametrizing families for $n$-dimensional smooth Fano varieties. Then the corollary is immediate from Theorem 1.2. ∎
We finish this section, by proving that the analogous statement to Corollary 1.3 fails if we replace cluster type with rational. The following example was kindly suggested to us by Brendan Hassett.
Theorem 5.3 (Theorem 1.4).
Let $n\geq 4$ be an integer. Let $f_{i}\colon\mathcal{X}_{i}\to T_{i}$ a finite set of smooth projective morphisms whose fibers are smooth rational Fano varieties of dimension $n$. Then, there exists an $n$-dimensional smooth rational Fano variety that is not isomorphic to a fiber of any $f_{i}$. That is, there do not exist finitely many families parametrizing $n$-dimensional smooth rational Fano varieties.
Proof.
Let $B$ be the open subset of ${\mathbb{P}}^{59}$ parametrizing all smooth $(2,2)$ hypersurfaces in ${\mathbb{P}}^{2}\times{\mathbb{P}}^{3}$. Let $U\rightarrow B$ be the universal family, and let $\pi\colon\mathbb{P}^{n-4}\times U\to B$ be the composition with the projection to $U$. By [28], we know that the set $B_{\rm rat}$ parametrizing rational fibers of $\pi$ is a countable union of proper closed subset of $B$, and $B_{\rm rat}\subset B$ is Zariski (and even Euclidean) dense. We argue that there is a closed point $b\in B_{\rm rat}$ for which $\mathbb{P}^{n-4}\times U_{b}$ is not isomorphic to a fiber of any $f_{i}$.
Let $\mathcal{H}_{\mathbb{P}^{11}}$ be the Hilbert scheme parametrizing closed subschemes of $\mathbb{P}^{11}$ with the same Hilbert polynomial as a $(2,2)$ hypersurface in $\mathbb{P}^{2}\times\mathbb{P}^{3}$ under the Segre embedding, and let $\mathcal{U}_{\mathbb{P}^{11}}\to\mathcal{H}_{\mathbb{P}^{11}}$ be the universal family. Let $\mathcal{H}_{\mathbb{P}^{2}\times\mathbb{P}^{3}}\subset\mathcal{H}_{\mathbb{P}^{11}}$ be the subset parametrizing such subschemes that are contained in $\mathbb{P}^{2}\times\mathbb{P}^{3}$. Then $\mathcal{H}_{\mathbb{P}^{2}\times\mathbb{P}^{3}}$ parametrizes $(2,2)$ hypersurfaces in $\mathbb{P}^{2}\times\mathbb{P}^{3}$, and by construction, $\mathcal{H}_{\mathbb{P}^{2}\times\mathbb{P}^{3}}\subset\mathcal{H}_{\mathbb{P}^{11}}$ is a locally closed subset. The family $\mathbb{P}^{n-4}\times\mathcal{U}_{\mathbb{P}^{11}}\to\mathcal{H}_{\mathbb{P}^{11}}$ induces a morphism $\mathcal{H}_{\mathbb{P}^{11}}\to\mathcal{H}_{\mathbb{P}^{12(n-4)+11}}$ to the corresponding component of the Hilbert scheme of $\mathbb{P}^{12(n-4)+11}$. Let $\mathcal{H}_{0}\subset\mathcal{H}_{\mathbb{P}^{12(n-4)+11}}$ denote the image of $\mathcal{H}_{\mathbb{P}^{2}\times\mathbb{P}^{3}}$; this is a constructible subset. Let $\phi_{B}\colon B\to\mathcal{H}_{\mathbb{P}^{2}\times\mathbb{P}^{3}}$ be the induced morphism to the Hilbert scheme. The image $\phi_{B}(B_{\rm rat})\subset\phi_{B}(B)$ is a proper subset that is Zariski dense. Furthermore, if $\psi_{B}\colon B\to\mathcal{H}_{\mathbb{P}^{2}\times\mathbb{P}^{3}}\to\mathcal{H}_{0}$ is the composition, then $\psi_{B}(B_{\rm rat})\subset\psi_{B}(B)=\mathcal{H}_{0}$ is again a Zariski dense proper subset.
Now we consider the families $f_{i}\colon\mathcal{X}_{i}\to T_{i}$. Since embedding dimension is upper semicontinuous in families, we may replace the $T_{i}$ by a (Zariski) locally closed stratification so that for each $i$, the fibers of $f_{i}$ all have embedding dimension $\leq 12(n-4)+11$ or all have strictly higher embedding dimension. Let $i$ be an index for which this embedding dimension is $\leq 12(n-4)+11$. Then we get an induced morphism $\psi_{i}\colon T_{i}\to\mathcal{H}_{\mathbb{P}^{12(n-4)+11}}$ to the Hilbert scheme, and the image $\psi_{i}(T_{i})$ is constructible. Then the intersection $\psi_{i}(T_{i})\cap\mathcal{H}_{0}$ is a constructible subset of $\mathcal{H}_{0}$, so it is a finite union of locally closed subsets $W_{ij}$ of $\mathcal{H}_{0}$. Each $W_{ij}$ must be contained in a proper Zariski closed subset of $\mathcal{H}_{0}$; otherwise, $W_{ij}$ would contain a nonempty open subset of $\mathcal{H}_{0}$ and hence would intersect the complement of $\psi_{B}(B_{\rm rat})$, implying that $f_{i}$ has an irrational fiber. Therefore, the intersection $\psi_{i}(T_{i})\cap\mathcal{H}_{0}$ is not dense in $\mathcal{H}_{0}$. We conclude that there exists a closed point $b\in\psi_{B}(B_{\rm rat})$ whose image in $\mathcal{H}_{\mathbb{P}^{12(n-4)+11}}$ is not contained in $\bigcup_{i}\psi_{i}(T_{i})$; that is, no fiber of any $f_{i}$ is isomorphic to the rational variety $\mathbb{P}^{n-4}\times U_{b}$. ∎